# Taylor-Fourier integration

### v1

@@ Abstract
In this paper we introduce an algorithm which provides approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions that after an appropriate change of variables can be written as a non-autonomous system with $(2\pi/\omega)$-periodic dependence on $t$. The proposed approximate solutions are written in closed form as functions $X(t,\omega\,t)$ where $X(t,\theta)$ is, (i) a truncated Fourier series in $\theta$ for fixed $t$, and (ii) a truncated Taylor series in $t$ for fixed $\theta$ (that is the reason for the name of the proposed integrators). Such approximations are intended to be uniformly accurate in $\omega$ (in the sense that their accuracy is not deteriorated as $\omega\to\infty$). This feature implies that Taylor-Fourier approximations become more efficient than the application of standard numerical integrators for sufficiently high basic frequency $\omega$. The main goal of the paper is to propose a procedure to efficiently compute such approximations by combining power series arithmetic techniques and the FFT algorithm. We present numerical experiments that demonstrate the effectiveness of our approximation method through its application to well-known problems of interest.
Keywords: Taylor series, Fourier methods, oscillatory problems, numerical integrators.
@@ 1 Introduction
In this paper we consider differential systems of the form
@@
where ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a smooth map, $\omega$ is a positive real parameter, and $A$ is a real $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$. Hence, the matrix exponential $e^{\theta A}$ is $(2\pi)$-periodic in $\theta$ and, therefore, the solution ${\bf x}(t)=e^{t\omega A}{\bf x}(0)$ of (1) with ${\bf g}\equiv{\bf 0}$ is $(2\pi/\omega)$-periodic in $t$. If $\omega$ is large, the solution of the homogeneous problem associated to (1) is highly oscillatory. Examples of differential systems (1) with these properties can be found, for instance, when dealing with semi-discretized dispersive partial differential equations [3, 9, 17], or in some formulations of the equations of motion in satellite dynamics [16].
We focus on approximating the solution ${\bf x}(t)$ of (1) supplemented with the initial condition
@@
$\frac{d}{dt}{\bf y}={\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf x}_{0},$ (4)
@@
which is $(2\pi)$-periodic in $\theta$.
Provided that the basic frequency $\omega$ is high enough, solving numerically (4) by a Runge-Kutta method demonstrates superior efficiency compared to the direct application of the same Runge-Kutta scheme to the original equation (1).
@@
From now on we will focus on computing approximations to the solution ${\bf y}(t)$ of the initial value problem (4) which are valid on an appropriate interval around $t=0$. In the highly oscillatory case, for high enough basic frequency $\omega$, the larger is the basic frequency $\omega$, the more computational effort will be required by the Runge-Kutta method to achieve a prescribed accuracy. In the present work, we aim at obtaining a closed form approximation to the solution ${\bf y}(t)$ of (4) that exhibits uniform accuracy across varying frequencies $\omega$. This implies that comparable accuracy is attained with the same computational effort even for arbitrarily high frequencies. The Taylor-Fourier approximations that we propose here draw inspiration from previous works on the numerical integration of oscillatory problems (see e.g. [4], [7], and references therein), and are related to the methods proposed in [5] and in [2].
For a prescribed positive integer $M$, we propose to compute a sequence
$\left\{{\bf y}^{[1]}(t),{\bf y}^{[2]}(t),\ldots,{\bf y}^{[d]}(t),\ldots\right\}$
of approximations to the solution ${\bf y}(t)$ of (4) of the form
${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right),\quad d=1,2,\ldots,$ (6)
with appropriate coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, and starting from ${\bf y}^{[0]}(t)\equiv{\bf x}_{0}$ (this means that we can use (6) with $d=0$, setting ${\bf y}_{0,0}^{[0]}={\bf x}_{0}$). We will refer to (6) as the (M,d)-Taylor-Fourier approximation to the solution of (4).
In order to compute the successive approximations (6) to the solution of (4), for $d=0,1,\ldots$, we do the following:
  • Given ${\bf y}^{[d]}(t)$, we first insert it into the right hand side of (4), and then we determine an approximation ${\bf z}^{[d]}(t)$ to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ of the form
    ${\bf z}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\left(\sum_{j=0}^{d}t^{j}\,{\bf z}_{k,j}^{[d]}\right).$ (7)
    This approximation, to be defined precisely in Section 2, is determined by combining truncated Taylor expansions and trigonometric interpolation of periodic functions. That is the reason for the name of the proposed integrators.
  • Once ${\bf z}^{[d]}(t)$ has been obtained, we use it as the right-hand side of (4) and integrate to obtain
@@
    It is straightforward to check that, by construction, ${\bf y}_{k,d+1}^{[d+1]}=0$ for $k\neq 0$.
@@
The aim of the present paper is: (i) to introduce and give a precise definition of Taylor-Fourier approximations, (ii) to propose a procedure to compute them efficiently by making extensive use of the FFT algorithm, and (iii) to illustrate the application of the proposed approximation method to certain well known problems of interest. Our findings suggest that, provided that the basic frequency $\omega$ is high enough, Taylor-Fourier approximations can achieve comparable precision to standard numerical integration methods applied to the transformed system (4), while demanding less CPU time.
We want to stress that employing Taylor-Fourier approximations may be of interest even in cases where the application of standard numerical integrators to the transformed system (4) proves to be very efficient. Indeed, the algorithm introduced in the present paper provides, for given $d$ and $M$, closed expressions
${\bf X}(t)=e^{t\omega A}\left(\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right)$ (9)
of approximate solutions of (1)-(2), valid for a relatively large time interval and that, once computed, can be efficiently evaluated at arbitrary times $t$. This is in contrast with most time integrators commonly used, which generate approximations to the exact solution of the problem either at a prescribed final time $T$ or at a finite set $0<t_{1}<t_{2}<\cdots\leq T$ of intermediate times. (In the later case, approximations to the solutions at arbitrary times $t\in[0,T]$ are typically obtained with piecewise interpolating polynomials.)
In our numerical experiments, we will compare the $(M,d)$-Taylor-Fourier approximations (6) to the solution ${\bf y}(t)$ of (4) (equivalently, the approximation (9) to the solution ${\bf x}(t)$ of (1)–(2)) with the approximations obtained by numerically integrating (4) with highly efficient Runge-Kutta methods (equivalently, numerically integrating (1)–(2) with Lawson’s generalized Runge-Kutta methods). A state-of-the-art highly performant library for the numerical integration of differential equations, DifferentialEquation.jl [15], implemented in julia language [1], is freely available and can be used to integrate the transformed system (4) with optimized adaptive explicit Runge-Kutta integrators. In order to compare such numerical solvers with the integrators proposed in this paper, we have carried out a careful implementation of the Taylor-Fourier methods in julia language, paying special attention to computational aspects such as memory management and efficient power series arithmetic.
@@
In Section 2 we give a precise definition of $(M,d)$-Taylor-Fourier approximations, and demonstrate their applicability to problem classes extending beyond real semi-linear problems of the form (1). Section 3 is devoted to describe an algorithm designed to efficiently compute Taylor-Fourier approximations. In Section 4 we present a couple of relevant numerical examples where the performance of Taylor-Fourier integrators is analyzed. Finally, in Section 5, we summarize our conclusions and outline avenues for future research on Taylor-Fourier methods.
2 Taylor-Fourier methods
2.1 Definition of Taylor-Fourier approximations
So far, we have not specified how to determine the approximation (7) to ${\bf f}(\omega t,{\bf y}^{[d]}(t))$ in the procedure introduced in Section 1. We first note that (6) and (7) can be rewritten as
${\bf y}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\omega t),\qquad{\bf z}^{[d]}(t)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\omega t),$
@@ 2.1 Definition of Taylor-Fourier approximations
respectively, where each component of the $D$-vector functions ${\bf Y}_{j}^{[d]}(\theta)$ and ${\bf Z}_{j}^{[d]}(\theta)$ is a trigonometric polynomial of degree $M$ in the angular variable $\theta$, that is,
${\bf Y}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf y}_{k,j}^{[d]},\qquad{\bf Z}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j}^{[d]},$
with coefficients ${\bf y}_{k,j}^{[d]},{\bf z}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$.
Given a positive integer $M$ and ${\bf Y}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$, we will determine ${\bf Z}_{j}^{[d]}(\theta)$, $j=0,1,\ldots,d$ by imposing the following conditions:
  • For each $\theta_{n}=n\pi/M$, $n=0,1,2,\ldots,2M-1$,
    ${\bf f}\left(\theta_{n},\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta_{n})\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta_{n})+\mathcal{O}(t^{d+1}){\color[rgb]{1,0,0},}$ (10)
    as $t\to 0$.
  • For $j=0,1,\ldots,d$,
    ${\bf z}_{-M,j}^{[d]}={\bf z}_{M,j}^{[d]}.$ (11)
The first condition is aimed at ensuring that
$\sum_{j=0}^{d}t^{j}\,{\bf Z}_{j}^{[d]}(\theta)\approx{\bf f}\left(\theta,\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)\right)$
for all $\theta\in[0,2\pi]$ and for sufficiently small $t\in\mathbb{R}$, whereas the second one is included to ensure uniqueness.
To sum up, given a positive integer $M$, the successive approximations (6) to the solution of (4) are uniquely determined by (10)–(11) and (8). We will refer to the approximation (6) as the $(M,d)$-Taylor-Fourier approximation.
2.2 Application to more general equations
Going back to equations in the real domain, Taylor-Fourier methods are not limited to semi-linear equations of the form (1). They can be applied more generally to problems of the form
@@ 2.2 Application to more general equations
    $\frac{\partial}{\partial t}\varphi_{t}({\bf x})={\bf r}(\varphi_{t}({\bf x})){\color[rgb]{1,0,0}.}$ (13)
    In that case, problem (12) is transformed with the change of variables
    ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$
    into (4), where
@@
${\bf x}^{[d]}(t)=\varphi_{\omega\,t}\left({\bf y}^{[d]}(t)\right),$ (15)
where
${\bf y}^{[d]}(t)=\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}$
is the $(M,d)$-Taylor-Fourier approximation to the solution of (4)–(14).
$\begin{split}\frac{d}{dt}\phi_{j}&=k_{j}\omega+g_{j}({\bf x}),\quad j=1,\ldots,d,\\ \frac{d}{dt}a_{j}&=g_{d+j}({\bf x}),\quad j=1,\ldots,D-d,\\ \end{split}$ (16)
where ${\bf x}=(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})$, $k_{j}\in\mathbb{Z}$, and for each $j$, $g_{j}$ is $2\pi$-periodic in each of the angle variables $\phi_{j}$. In that case,
$\varphi_{t}(\phi_{1},\ldots,\phi_{d},a_{1},\ldots,a_{D-d})=\left(\begin{matrix}\phi_{1}+k_{1}\,t\\ \vdots\\ \phi_{d}+k_{d}\,t\\ a_{1}\\ \vdots\\ a_{D-d}\\ \end{matrix}\right),$
and the change of variables ${\bf x}(t)=\varphi_{t\omega}({\bf y}(t))$, with ${\bf y}=(\tilde{\phi}_{1},\ldots,\tilde{\phi}_{d},\tilde{a}_{1},\ldots,\tilde{a}_{D-d})$, transforms the system (16) into
$\frac{d}{dt}\tilde{\phi}_{j}$ $=g_{j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,d,$
$\frac{d}{dt}\tilde{a}_{j}$ $=g_{d+j}(\tilde{\phi}_{1}+k_{1}\,\omega\,t,\ldots,\tilde{\phi}_{d}+k_{d}\,\omega\,t,\tilde{a}_{1},\ldots,\tilde{a}_{D-d}),\quad j=1,\ldots,D-d.$
3 Efficient implementation of Taylor-Fourier integrators
We will next show, for given positive integers $M$ and $d$, and given vector coefficients ${\bf y}_{k,j}^{[d]}\in\mathbb{C}^{D}$, $0\leq j\leq d$, $-M\leq k\leq M$, how to compute efficiently the new coefficients ${\bf y}_{k,j}^{[d+1]}\in\mathbb{C}^{D}$, $10\leq j\leq d+$ 1, $-M\leq k\leq M$. This computation is carried out in four sequential steps (for simplicity of notation, we omit superscripts $[d]$ in ${\bf Y}_{j}^{[d]}$ and ${\bf Z}_{j}^{[d]}$, when they are not necessary). We recall that $\theta_{n}=n\pi/M,\quad n=0,1,\ldots,2M-1$.

### v2

# Taylor-Fourier approximation

@@ Abstract
In this paper, we introduce an algorithm that provides approximate solutions to semi-linear ordinary differential equations with highly oscillatory solutions, which, after an appropriate change of variables, can be rewritten as non-autonomous systems with a $(2\pi/\omega)$-periodic dependence on $t$. The proposed approximate solutions are given in closed form as functions $X(\omega t,t)$, where $X(\theta,t)$ is (i) a truncated Fourier series in $\theta$ for fixed $t$ and (ii) a truncated Taylor series in $t$ for fixed $\theta$, which motivates the name of the method.
These approximations are uniformly accurate in $\omega$, meaning that their accuracy does not degrade as $\omega\to\infty$. In addition, Taylor-Fourier approximations enable the computation of high-order averaging equations for the original semi-linear system, as well as related maps that are particularly useful in the highly oscillatory regime (i.e., for sufficiently large $\omega$).
The main goal of this paper is to develop an efficient procedure for computing such approximations by combining truncated power series arithmetic with the Fast Fourier Transform (FFT). We present numerical experiments that illustrate the effectiveness of the proposed method, including applications to the nonlinear Schrödinger equation with non-smooth initial data and a perturbed Kepler problem from satellite orbit dynamics.
Keywords: Taylor series, Fourier methods, oscillatory ODE problems, numerical approximations.
@@ 1 Introduction
In this paper we are interested in systems of ordinary differential equations of the form
$\frac{d}{dt}{\bf x}=\mathcal{A}{\bf x}+{\bf g}({\bf x}),$ (1)
where ${\bf g}:\mathbb{R}^{D}\to\mathbb{R}^{D}$ is a smooth map and $\mathcal{A}$ is a real $D\times D$ matrix with purely imaginary eigenvalues.
Although the core concepts and techniques can be naturally extended to the general case of arbitrary purely imaginary eigenvalues, we primarily focus on the fully resonant case, in which the ratio of any pair of eigenvalues is rational. In this scenario, the solution ${\bf x}(t)=e^{t\mathcal{A}}{\bf x}(0)$ to (1) with ${\bf g}\equiv{\bf 0}$ becomes $(2\pi/\omega)$-periodic in $t$, where $\omega$ is the fundamental frequency (i.e., all eigenvalues of $\mathcal{A}$ are integer multiples of $\omega i$.)
We thus assume that $\mathcal{A}=\omega\,A$, where $A$ is a real $D\times D$ matrix whose eigenvalues are integer multiples of the imaginary unit $i$. We focus on approximating the solution ${\bf x}(t)$ of
@@
supplemented with the initial condition
@@
Examples of differential systems (2) with these properties can be found, for instance, when dealing with semi-discretized dispersive partial differential equations [3, 11, 19], or in some formulations of the equations of motion in satellite dynamics [18].
In the present work, we aim at obtaining a closed form approximation to the solution ${\bf x}(t)$ of (2) that exhibits uniform accuracy across varying frequencies $\omega$ for a time-interval (centered at $t=0$) of length independent of $\omega$.
@@
$\frac{d}{dt}{\bf y}={\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf y}_{0},$ (5)
@@
which is $(2\pi)$-periodic in $\theta$, and ${\bf y}_{0}={\bf x}_{0}$.
@@
The Taylor-Fourier approximations that we propose here draw inspiration from previous works on the analysis of numerical integration of oscillatory problems:
  • The technique of modulated Fourier expansions  (see e.g. [8, 10], and references therein), when applied to (5) would lead to a formal expansion
    $\sum_{k\in\mathbb{Z}}e^{ik\omega t}{\bf y}_{k}(t)$ (7)
    of the solution ${\bf y}(t)$ of (5), where ${\bf y}_{k}(t)$ ($k\in\mathbb{Z}$) are smoothly varying functions (i.e., their derivatives are bounded independently of $\omega$). An approximation of ${\bf y}(t)$ can be obtained from (7) by replacing each ${\bf y}_{k}(t)$ by its $d$th degree truncated Taylor expansion centered at $t=0$, and by discarding the terms corresponding to Fourier modes with $|k|>M$, that is,
    ${\bf y}(t)\approx\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}{\bf y}_{k,j}.$ (8)
    The $(M,d)$-Taylor-Fourier approximations that we propose in the present paper are precisely of the form (8).
  • In [5], the high-order averaging of systems of the form (5) is studied by expanding its solution ${\bf y}(t)$ as a formal series indexed by multi-indices $(k_{1}\cdots k_{\ell})\in\mathbb{Z}^{\ell}$. Such series expansion can be best described by parametrizing the equations (5) with a dummy variable $\epsilon$ as
    $\frac{d}{dt}{\bf y}=\epsilon\,{\bf f}(\omega t,{\bf y}),\quad{\bf y}(0)={\bf y}_{0}.$ (9)
    Its solution ${\bf y}(t)$ is expanded as
    ${\bf y}(t)={\bf y}_{0}+\sum_{\ell=1}^{\infty}\epsilon^{\ell}\sum_{(k_{1}\cdots k_{\ell})\in\mathbb{Z}^{\ell}}\alpha_{k_{1}\cdots k_{\ell}}(t)\,{\bf f}_{k_{1}\cdots k_{\ell}}({\bf y}_{0}),$
    where each ${\bf f}_{k_{1}\cdots k_{\ell}}$ is a map that is independent of $\omega$, and each $\alpha_{k_{1}\cdots k_{\ell}}(t)$ is a scalar function that only depends on $t$ and $\omega$ and the multi-index $(k_{1},\ldots,k_{\ell})$, but is independent from the actual map ${\bf f}(\theta,{\bf y})$. More precisely, each $\alpha_{k_{1}\cdots k_{\ell}}(t)$ is a linear combination with rational coefficients of terms of the form $\omega^{j-\ell}t^{j}e^{i\,k\,\omega\,t}$, $0\leq j\leq\ell$. An equivalent expansion with series indexed by mode-colored rooted trees was derived in [4]. After collecting terms with equal powers of $\epsilon$ such expansions of the solution ${\bf y}(t)$ can be written as
    ${\bf y}(t)={\bf y}_{0}+\sum_{\ell=1}^{\infty}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf y}_{k,\ell,j}.$ (10)
    As shown in [5, Theorem 3.1], under suitable conditions on the map ${\bf f}(\theta,{\bf y})$, there exists $\eta>0$ independent of $\omega$ such that the series (10) is absolutely and uniformly convergent for $|\epsilon\,t|\leq\eta$. An essential insight that inspired the consideration of Taylor-Fourier approximations, as presented in this paper, is the realization [4] that the approximations derived from the expansion (10) by truncating terms involving higher powers of $\epsilon$,
    ${\bf y}^{[d]}(t)={\bf y}_{0}+\sum_{\ell=1}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf y}_{k,\ell,j},$ (11)
    can be systematically constructed by combining Picard iterations with the truncation of the series expansion in powers of $\epsilon$. Indeed, starting with ${\bf y}^{[0]}(t)={\bf y}_{0}$, for $d=1,2,3,\ldots$, one first determines
    ${\bf z}^{[d]}(t)=\sum_{\ell=0}^{d}\epsilon^{\ell}\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{\ell}t^{j}{\bf z}_{k,\ell,j}$ (12)
    uniquely such that
    ${\bf f}(\omega t,{\bf y}^{[d]}(t))={\bf z}^{[d]}(t)+\mathcal{O}(\epsilon^{d+1}),$ (13)
    and then obtains
@@
    In the particular case where $\epsilon=1$, one obtains a series expansion of the solution ${\bf y}(t)$ of (5) that is absolutely and uniformly convergent for $|t|\leq\eta$, where $\eta>0$ is independent of $\omega$.
    Taylor-Fourier approximations, as presented in this paper, emerged as a practical computational method for approximating such expansions. To construct approximations of the form (8), an iterative process similar to (12)–(14) is employed. However, instead of truncating the series in powers of $\epsilon$, truncation is applied to powers of $t$, and the exact Fourier series expansion is replaced with trigonometric interpolation.
Our Taylor-Fourier approximations are also related to the methods proposed in [2] for a more restricted class of problems of the form (2).
@@
The aim of the present paper is: (i) to introduce and give a precise definition of Taylor-Fourier approximations, (ii) to relate them to high order stroboscopic averaging, (iii) to propose a procedure to compute them efficiently by making use of truncated power series arithmetic and the FFT algorithm, and (iv) to illustrate the application of the proposed approximation method to certain well known problems of interest.
For given $M$ and $d$, the algorithm introduced in this paper provides closed-form expressions
${\bf x}(t)\approx e^{t\omega A}\left(\sum_{k=-M}^{M}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}\right)$ (15)
of approximate solutions to (2)-(3). These approximations maintain uniform accuracy in $\omega$ and remain valid over a time interval independent of $\omega$. Once computed, these approximations can be efficiently evaluated at arbitrary times $t$ without additional costly computations. This contrasts with standard numerical integration methods, which produce approximations to the exact solution at a finite sequence of intermediate times $0<t_{1}<t_{2}<\cdots\leq T$, and whose accuracy typically deteriorates for large values of $\omega$.
The Taylor-Fourier approximations presented in this work are not intended to compete with numerical integration methods in terms of computational efficiency. Rather, their primary purpose is to provide reliable reference solutions over large time intervals. Additionally, they can serve as a valuable tool for assessing the accuracy and effectiveness of methods that approximate the solutions of high-order averaged equations.
We have carried out a careful implementation of the Taylor-Fourier methods in Julia language, paying special attention to computational aspects such as memory management and efficient truncated power series arithmetic.
@@
In Section 2 we give a precise definition of $(M,d)$-Taylor-Fourier approximations. Section 3 is devoted to describe an algorithm designed to efficiently compute truncated Taylor-Fourier series. In Section 4, we explore the relationship of the Taylor-Fourier technique with high-order stroboscopic averaging. We demonstrate in Section 5 the applicability of Taylor-Fourier approximations to problem classes extending beyond real semi-linear problems of the form (1) with fully resonant frequencies. In Section 6 we illustrate the application of the Taylor-Fourier algorithm to a couple of relevant examples. Finally, in Section 7, we summarize our conclusions.
2 Taylor-Fourier approximations
To lay the groundwork for the definition of fully discrete $(M,d)$-Taylor-Fourier approximations (8) to the solution ${\bf y}(t)$ of (5) for arbitrary pairs of positive integers $(M,d)$, we first introduce the semi-discrete $(\infty,d)$-Taylor-Fourier approximations, which correspond to the limit case as $M\to\infty$.
2.1 Semi-discrete Taylor-Fourier approximations
For each positive integer $d$, the $(\infty,d)$-Taylor-Fourier approximation to the solution ${\bf y}(t)$ of (5) is of the form
${\bf Y}^{[d]}(\omega t,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\omega t),$ (16)
@@ 2.1 Semi-discrete Taylor-Fourier approximations
where each component of the $D$-vector functions ${\bf Y}_{j}^{[d]}(\theta)$ is $(2\pi)$-periodic in the angular variable $\theta$. The function
${\bf Y}^{[d]}(\theta,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)$
is defined recursively. Starting with ${\bf Y}^{[0]}(\theta,t)={\bf y}_{0}$, ${\bf Y}^{[d+1]}(\theta,t)$ is determined for $d\geq 0$ as follows:
@@
    First, we determine the $d$th-order truncated expansion in powers of $t$ for ${\bf f}(\theta,{\bf Y}^{[d]}(\theta,t))$,
    ${\bf Z}^{[d]}(\theta,t)=\sum_{j=0}^{d}t^{j}{\bf Z}_{j}^{[d]}(\theta)={\bf f}\left(\theta,\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta)\right)+\mathcal{O}(t^{d+1}).$ (17)
@@
    Second, we determine ${\bf Y}^{[d]}(\theta,t)$ from the identity
[…]
Notice that if, for each $j$, the function ${\bf Y}_{j}^{[d]}(\theta)$ admits a Fourier expansion
${\bf Y}_{j}^{[d]}(\theta)=\sum_{k=-\infty}^{\infty}e^{ik\theta}{\bf y}^{[d]}_{k,j},$
then the $(\infty,d)$-Taylor-Fourier approximation
${\bf Y}^{[d]}(\omega t,t)=\sum_{k=-\infty}^{\infty}e^{ik\omega t}\sum_{j=0}^{d}t^{j}\,{\bf y}_{k,j}^{[d]}$ (19)
takes the form (8) in the limiting case $M\to\infty$.
2.2 Fully discrete Taylor-Fourier approximations
For the time being, let us assume that the Fourier expansion of ${\bf f}(\theta,{\bf y})$ contains only a finite number of nonzero modes. That is, there exists a positive integer $M_{0}$ such that
${\bf f}(\theta,{\bf y})=\sum_{k=-M_{0}}^{M_{0}}e^{ik\theta}\,\hat{\bf f}_{k}({\bf y}).$ (20)
In particular, if each component of ${\bf g}({\bf x})$ in (2) is a polynomial function of ${\bf x}\in\mathbb{R}^{D}$, then (6) admits such a finite Fourier expansion.
If condition (20) holds, then the semi-discrete Taylor-Fourier approximation (19) can also be expressed as a finite sum. More precisely, it can be shown that ${\bf y}^{[d]}_{k,j}=0$ for all $|k|>M=dM_{0}$.
In this case, the first step (17) in the recursive definition of $(\infty,d)$-Taylor-Fourier approximations can be split into two sub-steps. In the first sub-step, truncated expansions in powers of $t$ of univariate functions of $t$ are performed, while in the second one, we apply trigonometric interpolation for uniformly spaced nodes in $[0,2\pi]$. More precisely, we can consider $\theta_{n}=2n\pi/(2M+1)$, $n=0,1,2,\ldots,2M$, with $M\geq(d+1)M_{0}$, and
  • first compute for each node $\theta_{n}$ the vectors ${\bf Z}^{[d]}_{j}(\theta_{n})={\bf Z}_{n,j}^{[d]}\in\mathbb{C}^{D}$ from the truncated series expansion in powers of $t$
    ${\bf f}\left(\theta_{n},\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[d]}(\theta_{n})\right)=\sum_{j=0}^{d}t^{j}\,{\bf Z}_{n,j}^{[d]}+\mathcal{O}(t^{d+1})$ (21)
  • and then, for each $j=0,1,2,\ldots,d$, uniquely determine the functions
    ${\bf Z}_{j}^{[d]}(\theta)=\sum_{k=-M}^{M}e^{ik\theta}{\bf z}_{k,j}^{[d]}$ (22)
    from the conditions
    ${\bf Z}_{j}^{[d]}(\theta_{n})={\bf Z}_{n,j}^{[d]},\quad\theta_{n}=\frac{2n\pi}{2M+1},\quad n=0,1,2,\ldots,2M.$
The expansion (21) can be computed by evaluating the left-hand side of (21) with truncated power series arithmetic. As for the trigonometric interpolation, it is efficiently performed with the application of the Discrete Fourier Transform (DFT) (see Section 3). Since the efficiency of DFT-based trigonometric interpolation is optimal for number of nodes that are powers of 2, interpolation with an even number of nodes is preferred in practice. We thus replace the interpolatory conditions (having an odd number of nodes) above by the alternative conditions with an even number of nodes,
${\bf Z}_{j}^{[d]}(\theta_{n})={\bf Z}_{n,j}^{[d]},\quad\theta_{n}=\frac{n\pi}{M},\quad n=0,1,2,\ldots,2M-1,$ (23)
together with
${\bf z}_{-M,j}^{[d]}={\bf z}_{M,j}^{[d]},\quad j=0,1,\ldots,d.$ (24)
Condition (24) is imposed to ensure uniqueness of the trigonometric interpolation functions (22).
To sum up, (21)–(24) is a practical way of performing the first step (17) in the recursive definition of $(\infty,d)$-Taylor-Fourier approximations to the solution ${\bf y}(t)$ for (5) provided that the finiteness condition (20) and $M>(d+1)M_{0}$ hold. Otherwise, the functions ${\bf Z}^{[d]}_{j}(\theta)$ determined by (21)–(24) do not satisfy (17) exactly, as they will be affected by interpolation errors, that will depend on the actual value of the positive integer $M$. In that general scenario, the function
${\bf y}^{[M,d]}(t)={\bf Y}^{[M,d]}(\omega t,t)=\sum_{j=0}^{d}t^{j}\,{\bf Y}_{j}^{[M,d]}(\omega t),$ (25)
defined recursively by (21)–(24) and (18) (with the superscript $[d]$ replaced by $[M,d]$) can be seen as a fully discrete substitute of the $(\infty,d)$-Taylor-Fourier approximation. We will say that (25) is the $(M,d)$-Taylor-Fourier approximation to the solution ${\bf y}(t)$ for (5).
3 Efficient computation of Taylor-Fourier approximations
Next, we describe an efficient method to compute the new coefficients ${\bf y}_{k,j}^{[M,d+1]}\in\mathbb{C}^{D}$ for $10\leq j\leq d+$ 1 and $-M\leq k\leq M$, given the coefficients ${\bf y}_{k,j}^{[M,d]}\in\mathbb{C}^{D}$ for $0\leq j\leq d$ and $-M\leq k\leq M$. For simplicity, we omit the superscripts $[M,d]$ in the remainder of this section. This computation proceeds in four sequential steps.